Dr. Kingsley took 60 ounces of a 20% gold alloy and mixed it with a 30% gold alloy to get a 24% gold alloy. How many ounces of the 30% gold alloy did he use?
Answer by stanbon(75887)

Dr. Kingsley took 60 ounces of a 20% gold alloy and mixed it with a 30% gold alloy to get a 24% gold alloy. How many ounces of the 30% gold alloy did he use?
----------------------
20% DATA:
Amount = 60 oz ; amt. of gold = 0.20*60 = 12 oz
----------------------
30% DATA:
Amount = x oz ; amt. of gold = 0.30x oz
-----------------------
Alloy DATA:
Amount = x+60 oz ; amt of gold = 0.24(x+60) oz
-------------------------
EQUATION:
gold + gold = gold
12 + 0.30x = 0.24x + 14.4
0.06x = 2.4
x = 40 oz (Amount of 30% in the mixture)
===========================================
Cheers,
Stan H.